OpenStudy (anonymous):

what is an equation equivalent to y=4x-2

OpenStudy (anonymous):

what do you mean equivalent?

OpenStudy (anonymous):

idk it just says here

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yes 12x-4y=8

OpenStudy (anonymous):

5y-25=-10

OpenStudy (anonymous):

ok so then all you need to do is get y by itself and simplify

OpenStudy (anonymous):

4y-16x=10

OpenStudy (anonymous):

-x+4y=-8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you get your answer?

OpenStudy (anonymous):

no i think i jacked it up along the way

OpenStudy (anonymous):

ok so first one you get 4y=12x-8 divide by 4 to get y=3x-2

OpenStudy (anonymous):

is there an x somewhere in #2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where at?

OpenStudy (anonymous):

o my bad

OpenStudy (anonymous):

what do you mean in number 2

OpenStudy (anonymous):

third one 4y=16x+10 divide by 4 y=4x+(10/4)

OpenStudy (anonymous):

5y-25=-10

OpenStudy (anonymous):

25x

OpenStudy (anonymous):

-x+4y=-8 4y=x-8 y=(1/4)x-2

OpenStudy (anonymous):

5y=25x-10 y=5x-2

OpenStudy (anonymous):

so the third one y=4x+(10/4) is parallel if thats what its asking

OpenStudy (anonymous):

ok thx